OpenStudy (anonymous):

The three angles of a triangle are 3x, x + 10, and 2x - 40. Find the size of the smallest angle in the triangle?

OpenStudy (anonymous):

note that in a triangle, the sum of the internal angles add up to 180 degrees.

OpenStudy (anonymous):

the angles in a triangle add up to 180. so 3x+x+10+2x-40=180 6x=210 x=210/6 x=35 The smallest angle in the triangle is 2x-40=2(35)-40=30

OpenStudy (anonymous):

ha thank you